OpenStudy (anonymous):

what is the sum of the coefficients of the expansion of (2x-1)^20? some answers 1 and some 0..what is the right answer?

OpenStudy (anonymous):

0 Refer to the solution attachment.

Directrix (directrix):

I think it is (2 - 1) ^ 20 = 1^20 = 1. I need to look up a theorem.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

:) welcome